Perform a multi-factor analysis for each week using both input factors together (labor & materials) Perform...

  1. Perform a multi-factor analysis for each week using both input factors together (labor & materials)
  2. Perform a single factor productivity analysis for each week using each of the two inputs separately (labor & materials)
DATA:          Input
Week Output Labor Amt Matls #
1 96 4 60.0
2 72 3 50.0
3 92 4 100.0
4 50 2 40.0
5 69 3 60.0
6 52 2 50.0
Materials = $50.00 per pound (#)
Labor = 8 hours/day @ 5 days/wk @ $25 per hour
Overhead = $1.50 per labor hour
Answer #1

We have ,

Multifactor productivity for each week=Output/Total cost of inputs

=Output/(Labor cost+Material cost+ overhead cost)

Labor cost of a week =no of workers in a week * 40* 25

Material cost of a week=Material *50

Overhead cost =no of workers in a week * 40* 1.50

Single factory productivity :

Labor productivity =Total output /Total labor hours

Labor productivity of week 1=F3/(D3*40)=0.60 units per labor hour

Similarly for rest of the weeks.

Material productivity =Total output /Total material

Material productivity of week 1=F3/E3=1.60 units per pound of material.

Similarly for rest of the weeks.